OpenStudy (amistre64):

hope this formats :) \[\cfrac{d}{dx}\int_{0}^{x^4}cos(\sqrt{t})dt\]

OpenStudy (amistre64):

\[\cfrac{d}{dx}\int_{x^4}^{0}cos(\sqrt{t})dt\]

OpenStudy (zarkon):

\[\frac{d}{dx}\int\limits_{x^4}^{0}\cos(\sqrt{t})dt=-\frac{d}{dx}\int\limits_{0}^{x^4}\cos(\sqrt{t})dt=-\cos(\sqrt{x^4})4x^3\]

OpenStudy (anonymous):

yes flip and take derivative using chain rule and FTC

OpenStudy (anonymous):

get exactly what Zarkon wrote

OpenStudy (anonymous):

although you can probably do something with \[\sqrt{x^4}\]

OpenStudy (amistre64):

there zarkon is ... had to f5 it to get to see it :) I kept thining it was: \[2x\ cos(x^2)\]

OpenStudy (anonymous):

AMISTRE! :)

OpenStudy (amistre64):

howdy skittl

OpenStudy (amistre64):

\[\frac{d}{dx}\int_{a}^{b}F'(x)dx=F'(b)b'-F'(a)a'\] i just thought since it was cos(x^2) that a' was 2x :)

OpenStudy (anonymous):

no the "inside" function is x^4

OpenStudy (anonymous):

it is not \[\frac{d}{dx}\int _a^b\] it is \[\frac{d}{dx}\int_a^{g(x)}\]

OpenStudy (anonymous):

you are using ftc part in other form. the one you need says "derivative of integral is integrand"

OpenStudy (anonymous):

good morning!

OpenStudy (amistre64):

right, I had some wires crossed on the test I took in class :) i marked the right answer simply becasue it was tho only one that was close to my fauxpaux